What is the perimeter P of a rectangle whose length is 7 ft more than its width x
I have already tried to put in 2l(7)+2w

Answer by mananth(16946)

width =x
length = x+7
...
Perimeter = 2*(L+W)
P=2*(x+x+7)
P=2(2x+7)
...
m.ananth@hotmail.ca